(6) Solve the systen x-3y +2=1 2x - 4-22=2 X +24 -32= 1 RO
Answer #1

system Son: x-3y + z = 1 2x - y - 22=2 x+2y-32=1 AX=B (Non homogenous system) we woute [A ] where I argumented - 1 N 1 2 -1 -